OpenStudy (kaylak12345):

rosedewittbukater 4 years ago Graph y = 5x and y = log5x on a sheet of paper using the same set of axes. Use the graph to describe the domain and range of each function. Then identify the y-intercept of each function and any asymptotes of each function. (4 points)

OpenStudy (kaylak12345):

@triciaal

OpenStudy (kaylak12345):

@mathmate

OpenStudy (triciaal):

y-intercept is the value of y when x = 0

OpenStudy (triciaal):

domain set of x-values (input) controlled variable range set of output values from lowest to highest

OpenStudy (kaylak12345):

y=5^x isnt is no vertical asymptopes and all real domain values

OpenStudy (kaylak12345):

what about log one

OpenStudy (triciaal):

y = log5x do you mean base 5 or (5x)?

OpenStudy (kaylak12345):

base 5

OpenStudy (mathmate):

@kaylak12345 In the future, base 5 can be writen as log_5(x), the parentheses will separate the base from the independent variable. Even better, use LaTex to show \(log_5x\).

OpenStudy (kaylak12345):

oh ok

OpenStudy (kaylak12345):

i need help come back

OpenStudy (kaylak12345):

y=5^x=all real x-values and no vertical asymtopes,horizontal asymtope y=0, range all positive real numbers- y-intercept is 1, point is (0,1) y=logsub5x=all real y-values, what do i add

OpenStudy (kaylak12345):

@triciaal

OpenStudy (kaylak12345):

hello

OpenStudy (kaylak12345):

help

OpenStudy (kaylak12345):

nevermind i got it